Question 1 of 12 7 Points



What follows is a calculated question with 14 blanks.



If a=⟨ 4,7,8 ⟩ and b=⟨ 5,10,2⟩ then:




|a|= Blank 15. Calculate the answer by read surrounding text.




|b|= Blank 16. Calculate the answer by read surrounding text.




a⋅b= Blank 17. Calculate the answer by read surrounding text.




a×b=⟨ Blank 18. Calculate the answer by read surrounding text.



, Blank 19. Calculate the answer by read surrounding text.



, Blank 20. Calculate the answer by read surrounding text.






b×a=⟨ Blank 21. Calculate the answer by read surrounding text.



, Blank 22. Calculate the answer by read surrounding text.



, Blank 23. Calculate the answer by read surrounding text.






a×a=⟨ Blank 24. Calculate the answer by read surrounding text.



, Blank 25. Calculate the answer by read surrounding text.



, Blank 26. Calculate the answer by read surrounding text.






The angle between a and b is Blank 27. Calculate the answer by read surrounding text.



1
Expert's answer
2022-03-08T21:05:01-0500

Blank (15)


"\begin{array}{l}
a = \left\langle {4,\,7,\,8} \right\rangle & , & b = \left\langle {5,\,10,\,2} \right\rangle \\
\\
\left| a \right| = \sqrt {{{\left( 4 \right)}^2} + {{\left( 7 \right)}^2} + {{\left( 8 \right)}^2}} \\
\\
\left| a \right| = \sqrt {16 + 49 + 64} = 129
\end{array}\"



Blank (16)


"\left| b \right| = \sqrt {{{\left( 5 \right)}^2} + {{\left( {10} \right)}^2} + {{\left( 2 \right)}^2}} \\
\left| b \right| = \sqrt {25 + 100 + 4} = 129"



Blank (17)


"\begin{array}{l}
a = \left\langle {4,\,7,\,8} \right\rangle & , & b = \left\langle {5,\,10,\,2} \right\rangle \\
a\,\,b = \left\langle {4,\,7,\,8} \right\rangle \,\,\left\langle {5,\,10,\,2} \right\rangle \\
a\,\,b = \left( 4 \right)\left( 5 \right) + \left( 7 \right)\left( {10} \right) + \left( 8 \right)\left( 2 \right)\\
a\,\,b = 106
\end{array}\"



Blank (18, 19, 20)





Blank (21, 22, 23)



Blank (24, 25, 26)



Blank (27)


The angle between a and b is 


"\begin{array}{l}
a \cdot b = \left| a \right|\left| b \right|\,\cos \theta \\
\theta = {\cos ^{ - 1}}\left( {\frac{{a \cdot b}}{{\left| a \right|\left| b \right|}}} \right)\\
\theta = {\cos ^{ - 1}}\left( {\frac{{106}}{{\sqrt {129} \sqrt {129} }}} \right)\\
\theta = {34.74^0}
\end{array}\"